network outage this morning

| | Comments (2)
right now, my upstream believes the problem is not my equipment;  they gave me a 30 minute ETA, but it sounds like it's too early to tell what the problem is.   

As of this moment, we believe the problem is unrelated to the server move, but like I said, this is very early in the process and it's possible that's wrong.  I will update in 30.

update: 08:06  we appear to be back.   I will update when we know more about what the actual problem was.  

update: 8:25:  bowl and cerberos are still down, working on it...  bowl and cerberus are related to last nights move.  

update: 9:57

my upstream says:
There have been multiple links affected by this fiber cut,
It is still in progress of being repaired, however the primarly link to SVTIX
+should be working now.

We are still waiting for details on the cause of the disruption.

update at 2011-06-25:
" Sorry for the delay, the lates news we have is "I can tell you right now that
+there was a fiber cut due to construction excavating to put in a new water
+main. ", and we already request a RFO for this outage, and it take 7-10
+business day for them to process.


Thank you for dealing with this, I know it's early on that side of the world :)

Just note I can't access from before 14:00 BST today - it's getting to 2.5 hours downtime ignoring the planned maintenance. Please keep us updated!

Thanks, working now! - it appears the Xen Domain didn't exist when access was restored. According to syslog, system was off between 13:40:41 to 16:27:02 BST.

Leave a comment

About this Entry

This page contains a single entry by luke published on June 22, 2011 7:31 AM.

server move on Tuesday and Wednesday night was the previous entry in this blog.

lozenges crash is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.